OpenStudy (anonymous):

Solve: 27-3x/5x-45 -1 -3/5 -24/41 27/45

OpenStudy (anonymous):

-3/5

OpenStudy (anonymous):

-3/5

OpenStudy (anonymous):

So answer B

OpenStudy (anonymous):

haha thank you guys <3

OpenStudy (anonymous):

\[\frac{27-3x }{ 5x -45} \\ =\frac{ 3 ( 9-x) }{ 5(x -9)} \\ =\frac{ (-1)3 ( x-9) }{ 5(x -9)} \\ =\frac{ (-1)3 ( \cancel{x-9}) }{ 5 \cancel{x-9} } \\ =-\frac{ 3 }{5} \]